# Learning to combine top-down context and feed-forward representations under ambiguity with apical and basal dendrites

### v1

@@ Abstract
One of the most striking features of neocortical anatomy is the presence of extensive top-down projections into primary sensory areas. Notably, many of these top-down projections impinge on the distal apical dendrites of pyramidal neurons, where they exert a modulatory effect, altering the gain of responses. It is thought that these top-down projections carry contextual information that can help animals to resolve ambiguities in sensory data. However, it has yet to be demonstrated how such modulatory connections to the distal apical dendrites can serve this computational function. Here, we develop a computational model of pyramidal cells that integrates contextual information from top-down projections to apical compartments with sensory representations driven by bottom-up projections to basal compartments. When input stimuli are ambiguous and relevant contextual information is available, the apical feedback modulates the basal signals to recover unambiguous sensory representations. Importantly, when stimuli are unambiguous, contextual information which is irrelevant or opposes sensory evidence is appropriately ignored by the model. By generalizing the task to temporal sequences, we further show that our model can learn to integrate contextual information across time. Using layer-wise relevance propagation, we extract the importance of individual neurons to the prediction of each category, revealing that neurons that are most relevant for the overlap of categories receive the largest magnitude of top-down signals, and are necessary for solving the task. This work thus provides a proof-of-concept demonstrating how the top-down modulatory inputs to apical dendrites in sensory regions could be used by the cortex to handle the ambiguities that animals encounter in the real world.
@@ 1 Introduction
Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information [94, 88]. Contextual information can be derived from a number of sources, including the spatial and temporal domains [65], task demands [69], as well as other sensory modalities.
@@
Several key neuroanatomical and cellular features could support the computations associated with context integration. In the mammalian brain, context has been proposed to be conveyed by top-down feedback pathways which are abundant in sensory regions of the neocortex [66, 85, 72]. Perturbations of top-down pathways from higher order areas have been associated with delayed object recognition [75] as well as disrupted stimulus response curves in lower-order regions [104, 90], which may reflect deficits in contextual processing. Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images [68, 99] as well as context-invariant perception [91].
@@
In the neocortex, pyramidal neurons are thought to integrate feedforward and top-down, feedback information at their basal and apical dendrites, respectively [80, 101]. Notably, feedback has been proposed to modulate rather than drive neuronal activities [97]. This idea is supported by the unique morphological and physiological properties of pyramidal neurons, as activity in apical dendrites in vivo is generally not sufficient to drive somatic output responses alone [102, 79, 77, 73]. On the other hand, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity [78]. While the physiological properties of pyramidal neurons have been described extensively (see [101, ] for review), there is no consensus on the exact mechanism by which top-down signals from higher-order regions modulate somatic firing rates and update sensory representations.
Here, we propose a functional model that learns to use context only when such information is relevant, and ignores it otherwise, ultimately learning a useful mapping of top-down contextual information onto apical compartments. We examine character recognition tasks in cases where the characters are ambiguous and consider two different computational settings. First, the contextual representation is directly given to the model which learns to use this top-down input to resolve the ambiguities. Then, the representation of context is learned, from the temporal framing of the character in words or sequences of digits. In both settings, we show that the model learns to use apical modulation, driven by contextual inputs, to resolve sensory ambiguities. We find that a Hadamard (element-wise multiplicative) dendritic integration rule informed by physiological data enables strong performance in both settings of the task.
To gain insights into the learned mechanisms of top-down modulation involved in our model, we applied layer-wise relevance propagation [57] to identify the relative contributions of individual neurons and their apical compartments to network function. We identified a subset of neurons that are highly relevant for class-pairs which are selectively amplified by top-down signals arriving at their apical dendrites. We propose that, in principle, the representation of context can be learned independently from the impact that it has on ambiguity resolution. These findings provide a candidate mechanism for how neocortical pyramidal neurons integrate top-down contextual information at their apical dendrites to support robust perception in the face of ambiguous data.
@@ 2.1 Functional model of context integration in apical dendrites
To model how apical dendrites integrate context to refine perceptual representations, we developed a task where correct classification of ambiguous characters requires appropriate integration of contextual cues (Fig. 1a). In this task, we trained a generative model to construct a dataset of ambiguous handwritten digits and letters based on MNIST and EMNIST, respectively. This generative model synthesizes images that are ambiguous between pairs of classes (see Methods). The resulting ambiguity was empirically quantified with a linear classifier trained on unambiguous images (fig. 1b). Inspired by anatomical features of pyramidal neurons, we next developed a model that implements distinct dendritic compartments, a basal compartment for integrating sensory information and an apical compartment for integrating contextual information (Fig. 1c). In the model, a population of $n$ pyramidal neurons has two sources of inputs, represented as vectors: a set of basal inputs, $\mathbf{b}=[b_{1},\ldots,b_{n}]$, representing the sensory stream, and a set of apical inputs, $\mathbf{a}=[a_{1},\ldots,a_{n}]$, representing the top-down stream. Input images, $\mathbf{x}$, are encoded by a pretrained network $f$, with parameters $\Theta_{b}$. $f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pretrained). The output of $f(x)$ is the basal vector $\mathbf{b}$.
The firing rate of pyramidal neurons, $\mathbf{h}=[h_{1},\ldots,h_{n}]$, is then determined by a thresholded element-wise (neuron-wise) combination of the basal and apical input vectors:
@@
where $\odot$ is an element-wise multiplication and $\sigma$ is the rectified linear unit (ReLU) activation function. We refer to this integration rule as ’Hadamard’. As can be seen from equation 1, the impact of the apical input vector, $\mathbf{a}$, on the firing rate, $\mathbf{h}$, is determined by $\sigma(\mathbf{a})+1$, which has elements that are by design always greater than or equal to one. As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity. This modulated basal activity is ultimately the firing rate of the modelled neurons, in-line with previous experimental findings [105]. Under these constraints, apical inputs affect neuronal activity only under the condition that they exceed their activation threshold (when $a>0$), similar to the requirement for regenerative calcium spikes for apical inputs to influence somatic spiking in real pyramidal neurons [79]. The output of the model is read-out from a feedforward linear transformation of the pyramidal neurons firing rates:
@@
where
$\mathbf{U}$
is a pretrained linear transformation and $\mu$ represents a higher-order neuronal population. To obtain the apical inputs $\mathbf{a}$, the context, $\mathbf{c}$, (which is either derived from temporal context, or given directly, see below) is merged with the output and mapped onto the apical compartment using a multilayer perceptron (MLP), $g(\cdot)$, as follows:
@@
where $\oplus$ is the concatenation operation, $\sigma$ is the ReLU activation, and $g$ is a 2-layer MLP (1 hidden layer) with learnable parameters denoted by $\Theta_{a}$. The model must use the combination of context and input to determine the representation of both ambiguous and unambiguous characters. For ambiguous characters, which are a mixture of two categories, the readout of the representations of the model will yield two plausible interpretations (one corresponding to the target category, and one corresponding to the contradictory category). The contradictory category refers to the class that induces the most confounding readout classification. Categories other than the target or contradictory we refer to as irrelevant categories. To test whether the model can learn useful apical modulations, we trained the synaptic weights $\Theta_{a}$ with the objective $L$ of modulating basal representations to match the target latent representations under various scenarios of input and context by gradient descent; Fig. 1e; pseudocode 1, and 5) This setup ensures that the top-down model is trained on both unambiguous and ambiguous samples, and that the model must learn to provide the appropriate modulatory signal only when it is relevant, ensuring that our model could learn to ignore contextual information when appropriate. This implementation reflects the fact that, in the mammalian brain, top-down modulation is available from multiple sources and present regardless of whether the sensory inputs are ambiguous or not, and whether the contextual sources are helpful for resolving ambiguity or not.
@@
This suggests an approach where the problem of contextual integration is broken into two parts: (1) how apical dendrites locally learn to use contextual representations $c$ to solve the task and (2) how those contextual representations $c$ can be learned from the temporal frame.
@@
For the first part, our initial experiments assume that context, $c$, is given in the form of a one-hot vector, a scenario we refer to as the ’oracle context’ (see Methods). Subsequent experiments for the second part address the problem of learning a representation of context, $c$, taking inspiration from the neocortex, where context is, in part, represented by higher-order regions with broader windows of temporal integration [65] (Fig. 1c; see Methods).
@@
Figure 1: Functional model of context integration in apical dendrites. a, in our experimental paradigm, ambiguous characters (in yellow, top panels) are resolved by combining context signals (bottom panels). b, examples of unambiguous pairs and resulting ambiguous characters generated using EMNIST (left) and MNIST (right). c, model of pyramidal neurons with distinct dendritic compartments trained to solve a task of perceptual ambiguity. Before top-down modulation, two predictions have equal probability. After top-down modulation, context contributes to favoring one unambiguous representation. d, model implementation of dendritic integration using a hierarchical convolutional neural network. e, distinct combinations of input and context used to train the top-down network $g$. f, possible combinations of input and context used during training and analysis.
@@ 2.2 Resolving ambiguity in the oracle context
To first test if our model can categorize digits under contextual input represented as a one-hot vector (oracle context), we trained $g$ on the ambiguous MNIST/EMNIST task with a loss including all 5 scenarios of stimulus and context (see Methods). We assessed performance of the model in each scenario as the held-out test set accuracy on the pretrained category readout. and found model performance was significantly modulated by top-down signals (one-way ANOVA; $F_{4}$ = 10076.598; p $<$ 0.0001; n = 313 independent samples). Test set accuracy on ambiguous digits was $44.619\pm 8.073\%$ without top-down modulation, and became significantly higher with top-down modulation ($99.491\pm 1.306\%$; Tukey’s test; p $<$ 0.0001) if context was matching, but did not improve performance if context was irrelevant ($43.920\pm 8.160$; Tukey’s test; p = 0.4651; n = 313 independent replicates; Fig. 2a; Table 1).
[…]
@@
Figure 2: Ambiguity can be resolved only when integrating top-down signals. a, readout test set accuracy before and after combining top-down signals over training epochs for all input-context combinations. Dashed line corresponds to chance level. b, low dimensional projection (t-SNE) of $h$ for unambiguous inputs (top), ambiguous inputs before (bottom left), and after (bottom right) top-down modulation. c, silhouette scores for $h$, for each class pair and for unambiguous inputs (left), ambiguous inputs before (center), and after (right) top-down modulation. d, cosine distance between $h$ in unambiguous conditions and basal, apical compartments and firing rate in each scenario.
[…]
@@
Figure 3: Role of integration rules on ambiguity resolution. a, comparison between Hadamard and arithmetic sum integration rules. b, test set accuracy relative to contextual certainty using either integration rule in conditions where inputs are ambiguous and context is useful. c, same but in conditions where inputs are unambiguous and context contradicts the true label.

### v2

@@ Abstract
One of the hallmark features of neocortical anatomy is the presence of extensive top-down projections into primary sensory areas, with many impinging on the distal apical dendrites of pyramidal neurons. While it is known that they exert a modulatory effect, altering the gain of responses, their functional role remains an active area of research. It is hypothesized that these top-down projections carry contextual information that can help animals to resolve ambiguities in sensory data. One proposed mechanism of contextual integration is a non-linear integration of distinct input streams at apical and basal dendrites of pyramidal neurons. Computationally, however, it is yet to be demonstrated how such an architecture could leverage distinct compartments for flexible contextual integration and sensory processing when both sensory and context signals can be unreliable. Here, we implement an augmented deep neural network with distinct apical and basal compartments that integrates a) contextual information from top-down projections to apical compartments, and b) sensory representations driven by bottom-up projections to basal compartments. In addition, we develop a new multi-scenario contextual integration task using a generative image modeling approach. In addition to generalizing previous contextual integration tasks, it better captures the diversity of scenarios where neither contextual nor sensory information are fully reliable. To solve this task, this model successfully learns to select among integration strategies. Specifically, when input stimuli and contextual information are contradictory, the performance of deep neural networks augmented with our "apical prior" exceeds that of single-compartment networks with otherwise equivalent architecture. We further show that our model can learn to integrate contextual information across time by using a task with input sequences and a gated recurrent unit layer to generate contextual signals. Using layerwise relevance propagation, we extract the relevance of individual neurons to the model’s predictions, revealing that a sparse subset of neurons encoding features of the context-relevant categories receive the largest magnitude of top-down signals when context and sensory input are aligned. We further show that this sparse gain modulation is necessary for best performance on the task. Altogether, this suggests that the "apical prior" and the biophysically-inspired non-linear integration rule could be key components necessary for handling the ambiguities that animals encounter in the diverse contexts of the real world.
@@ 1 Introduction
Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information (Rao and Ballard, 1999; Mumford, 1992). Contextual information can be derived from a number of sources, including the spatial and temporal domains (Eichenbaum, 2017), task demands (Gilbert and Li, 2013), as well as other sensory modalities.
@@
Several key neuroanatomical and cellular features could support the computations associated with context integration. In the mammalian brain, context has been proposed to be conveyed by top-down feedback pathways which are abundant in sensory regions of the neocortex (Felleman and Van Essen, 1991; Markov et al., 2014; Harris et al., 2019). Perturbations of top-down pathways from higher order areas have been associated with delayed object recognition (Kar and DiCarlo, 2021) as well as disrupted stimulus response curves in lower-order regions (Wang et al., 2007; Nassi et al., 2013), which may reflect deficits in contextual processing. Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images (George et al., 2017; Spoerer et al., 2017) as well as context-invariant perception (Naumann et al., 2022), and context-dependent gating for multi-task learning Masse et al. (2018).
@@
In the neocortex, pyramidal neurons are thought to integrate feedforward and top-down, feedback information at their basal and apical dendrites, respectively (Larkum et al., 2007; Spruston, 2008). Notably, feedback has been proposed to modulate rather than drive neuronal activities (Sherman and Guillery, 1998). This idea is supported by the unique morphological and physiological properties of pyramidal neurons, as activity in apical dendrites in vivo is generally not sufficient to drive somatic output responses alone (Stuart and Spruston, 1998; Larkum et al., 1999; Larkum, 2013; Jarvis et al., 2018). However, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity (Larkum et al., 2004), serving as a potential stream that could carry contextual information. While the physiological properties of pyramidal neurons have been described extensively (see Spruston (2008) for review), there is no consensus on the exact mechanism by which top-down signals from higher-order regions modulate somatic firing rates and update sensory representations.
Here, we propose a deep neural network architecture in which neurons are augmented specifically with apical and basal compartments and whose computation respects known biophysical properties of pyramidal neurons. We design a training paradigm such that, according to the context, the model updates its higher-level representations of sensory inputs, via the interaction of basal and apical compartments, in which the latter carries top-down contextual information.
In tandem with the proposed framework, we develop a multi-scenario task with a dataset containing ambiguous images that are mixtures of two image categories, in which ambiguity can only be correctly resolved by integrating contextual information with sensory representations.
We consider two versions of the problem setting. The first simplifies the contextual representation, which is given directly via a top-down oracle, and the model only needs to learn the apical parameters to correctly update the somatic activity. In the second, more challenging one, the model must use the temporal sequence as context and learn the appropriate contextual representations, along with the apical parameters to update somatic activity and ultimately solve the task.
We show that the model learns to use apical modulation, driven by contextual inputs, to resolve sensory ambiguities. We find that a neuron-wise (element-wise) dendritic integration rule informed by physiological data enables strong performance in both settings of the task.
To gain insights into the learned mechanisms of top-down modulation, we applied a standard method from the explainable AI literature ((Bach et al., 2015)) to identify the relative contributions of individual neurons and their apical compartments to network function. We identified a subset of neurons that are highly relevant for the matching category, and selectivelyamplified by top-down signals arriving at their apical dendrites.
We propose that, in principle, 1) apical gain modulation of somatic activity and contextual integration are compatible, and is a simple yet efficient way to solve new tasks such as resolving perceptual ambiguity without altering what the model has previously learned, as it does not introduce or modify any synapses in the feed-forward pathway, and 2) that the apical gain modulatory signals and the representations of contextual information arriving at the aforementioned apical compartments can be decoupled. In a highly distributed system with local learning constraints such as the brain, the representation of context can be learned independently from the mapping of contextual signals to local populations to solve specific perceptual tasks.
These findings provide a candidate mechanism for how neocortical pyramidal neurons integrate top-down contextual information at their apical dendrites to support robust perception in the face of ambiguous data.
@@ 2.1 Functional model of context integration in apical dendrites
To model how apical dendrites integrate context to refine perceptual representations, we first developed a multi-scenario contextual integration task where correct classification of ambiguous characters requires appropriate integration of contextual cues (Fig. 1a). To create the task, we trained a generative model on handwritten digits and letters (MNIST and EMNIST, respectively) to construct a dataset with highly ambiguous characters. We used the generative model to synthesize images that are a mixture of two classes (see Methods). The ambiguity of the synthesized images was empirically quantified as the proportion of the confusion of a linear classifier trained on unambiguous images (fig. 1b). Inspired by anatomical features of pyramidal neurons, we next developed a model that implements distinct dendritic compartments, a basal compartment for integrating sensory information and an apical compartment for integrating contextual information (Fig. 1c). In the model, a population of $n$ pyramidal neurons has two sources of inputs, represented as vectors: a set of basal inputs, $\mathbf{b}=[b_{1},\ldots,b_{n}]$, representing the sensory stream, and a set of apical inputs, $\mathbf{a}=[a_{1},\ldots,a_{n}]$, representing the top-down stream. Input images, $\mathbf{x}$, are encoded by a pretrained network $f$, with parameters $\Theta_{b}$, where $f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pretrained). The output of $f(x)$ is the basal vector $\mathbf{b}$.
The firing rate of pyramidal neurons, $\mathbf{h}=[h_{1},\ldots,h_{n}]$, is then determined by a thresholded neuron-wise (element-wise) combination of the basal and apical input vectors:
@@
where $\odot$ is an neuron-wise multiplication and $\sigma$ is the rectified linear unit (ReLU) activation function implementing basal and apical non-linearities, respectively. As can be seen from equation 1, the impact of the apical input vector, $\mathbf{a}$, on the firing rate, $\mathbf{h}$, is determined by $\sigma(\mathbf{a})+1$. As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity, in-line with previous experimental reports of a multiplicative role for apical inputs (Waters et al., 2003). Under these constraints, apical inputs affect neuronal activity only when they exceed their activation threshold (when $a>0$), analogous to observations that apical inputs influence somatic spiking only if they evoke a dendritic calcium spike (Larkum et al., 1999). To monitor the ability of the model to solve the perceptual task, the output of the model is read-out from a linear transformation of the pyramidal neurons firing rates:
@@
where $\mu$ represents a higher-order neuronal population ($\mathbf{U}$ is a linear transformation to the latent space of the encoder $f$). To obtain the apical inputs $\mathbf{a}$, the context signal, $\mathbf{c}$, is concatenated with $\mu$, and mapped onto the apical compartment using a multilayer perceptron (MLP), $g(\cdot)$, as follows:
@@
where $\oplus$ is the concatenation operation, $\sigma$ is the ReLU activation, and $g$ is a MLP with 1 hidden layer, and its parameters are denoted by $\Theta_{a}$.
As a population, these pyramidal neurons integrate sensory input character representations onto their basal dendritic compartment on one hand, and contextual information onto their apical dendritic compartment on the other hand.
For ambiguous characters, which are a mixture of two categories, a readout of the representations of the model will yield two plausible interpretations, and contextual information is required to resolve the ambiguity. We consider the category which is compatible with the contextual input to be the target category. Conversely, a category which is a plausible interpretation of the input but not compatible with the contextual information is referred to as the contradictory category. We refer to all other categories as irrelevant categories.
To test whether the model can learn apical modulations that are helpful for resolving ambiguous stimuli, we trained the synaptic weights $\Theta_{a}$ with the objective of modulating basal representations to match the target $\mu$ (representation after linear transform). We optimize this objective simultaneously for a set of 5 input-context scenarios (unambiguous or ambiguous input; helpful, contradictory or irrelevant context (Fig. 1e) using a common gradient descent optimization algorithm (Adam, see pseudocode 1, and 5). The labels and corresponding descriptions we use for each scenario are provided explicitly below:
  • ambig.Match: ambiguous input, helpful (matching) context
  • unambig.Irrel: unambiguous input, irrelevant context
  • ambig.Irrel: ambiguous input, irrelevant context
  • unambig.Match: unambiguous input, helpful (matching) context
  • unambig.Contra: unambiguous input, contradictory context
This setup ensures that the top-down model is trained on both unambiguous and ambiguous examples, and that the model must learn to provide the appropriate modulatory signal only when contextual inputs provide useful information for solving the task, and ignore them when they are irrelevant or contradictory to sensory information. Our proposed implementation reflects the reality that, in the mammalian brain, top-down modulation is available from multiple sources regardless of the ambiguity of the sensory inputs or the relevance of the contextual information.
@@
This suggests an approach where the problem of contextual integration is broken into two parts: (1) how apical dendrites locally learn to use contextual representations $c$ to solve the multi-scenario task and (2) how useful contextual representations $c$ can be learned from available sources (e.g. temporal information).
@@
For the first part, our initial experiments assume that context, $c$, is given in the form of a one-hot vector, a scenario we refer to as the ’oracle context’ (see Methods). Subsequent experiments for the second part address the problem of learning a representation of context, $c$, taking inspiration from the neocortex, where context is, in part, represented by higher-order regions with broader windows of temporal integration (Eichenbaum, 2017) (Fig. 1c; see Methods).
@@
Figure 1: Functional model of context integration in apical dendrites. a, in our experimental paradigm, ambiguous characters (in yellow, top panels) are resolved by combining context signals (bottom panels). b, examples of unambiguous pairs and resulting ambiguous characters generated using EMNIST (left) and MNIST (right). c, model of pyramidal neurons with distinct dendritic compartments trained to solve a multi-scenario contextual integration task with ambiguity. Before top-down modulation, two predictions have equal probability. After top-down modulation, context contributes to favoring one unambiguous representation. d, implementation of dendritic integration in artificial neurons with apical and basal compartments. $f$ represents the feedforward mapping onto the basal compartment parameterized by the pretrained backbone, and $g$ the contextual top-down mapping onto the apical compartment, parameterized by its own neural network. The combination of the outputs $b$ and $a$ of $f$ and $g$, respectively, is applied neuron-wise via the multiplicative interaction $\sigma(b)*(1+\sigma(a))$. e, distinct combinations of input and context used to train the top-down network $g$. f, possible combinations of input and context used during training and analysis.
@@ 2.2 Resolving ambiguity in the oracle context
To first test if our model can solve our multi-scenario task under contextual input represented as a one-hot vector (oracle context), we trained $g$ on the ambiguous MNIST/EMNIST task with a loss function that includes a loss term for each scenario We assessed performance of the model in each scenario as the accuracy of the readout on the test set, and found significant impact of top-down modulation. (one-way ANOVA; $F_{4}$ = 10076.598; p $<$ 0.0001; n = 313 independent samples). Specifically, test set accuracy on ambiguous digits was $46.3\pm 0.3\%$ without top-down modulation, and became significantly higher with top-down modulation ($98.5\pm 0.3\%$; Tukey’s test; p $<$ 0.0001) if context was matching, but did not significantly alter performance if context was irrelevant ($43.920\pm 8.160$; Tukey’s test; p = 0.4651; n = 313 independent replicates; Fig. 2a; Table 2). Importantly, the model did not learn to rely solely on top-down context, since performance remained high when the model was given contradictory context (Fig. 2a; 2). To visualize the representations of the model on a population level, we projected unmodulated and modulated outputs $\mathbf{\mu}$ onto a two-dimensional manifold using t-SNE (Maaten (2008); Fig. 2b). Ambiguous images whose representations exhibited high overlap between class-pairs were effectively disentangled after integrating matching contextual inputs (paired t-test, t99 = 6.0293, p $<$ 0.0001 for silhouette scores between ambiguous images before and after integrating top-down inputs; Fig. 2c). These results show that our model learned to use contextual inputs provided to apical dendrites to modulate basal activity and resolve ambiguity.
@@
Figure 2: Our task is best solved when top-down signals are integrated correctly. a, readout test set accuracy before and after combining top-down signals over training epochs for all input-context combinations. Dashed line corresponds to chance level. b, low dimensional projection (t-SNE) of $h$ for unambiguous inputs (top), ambiguous inputs before (bottom left), and after (bottom right) top-down modulation. c, silhouette scores for $h$, for each class pair and for unambiguous inputs (left), ambiguous inputs before (center), and after (right) top-down modulation. d, cosine distance between $h$ in unambiguous conditions and basal, apical compartments and firing rate in each scenario.
Task Context signal ambig (Baseline) ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra aMNIST oracle (one-hot) $46.3\pm 0.3$ $98.5\pm 0.3$ $98.5\pm 0.1$ $97.9\pm 0.1$ $46.0\pm 0.8$ $96.0\pm 0.4$ aEMNIST oracle (one-hot) $46.5\pm 0.5$ $94.2\pm 1.4$ $98.9\pm 0.2$ $96.2\pm 0.4$ $47.1\pm 0.8$ $94.9\pm 0.5$
Table 1: Ambiguous MNIST and EMNIST (aMNIST, aEMNIST) test set classification accuracy across input-context scenarios with the hadamard integration rule. Results are expressed for each scenario as mean % $\pm$ standard deviation with 3 random seeds.
Next, we evaluated the importance of having topdown apical compartments to handle contextual inputs compared with only having basal compartments (akin to point-wise spiking neurons, (Table. 4). We found that, while there was not a large difference in the other 4 scenarios, there was an average increase in performance of 1.9% (MNIST) in the unambig.Contra scenario for the models which had apical compartments compared to those which only had basal compartments.
Next, focusing our analysis on MNIST, we evaluated the importance of the integration rule which combines basal and apical activities4, 3a).
We found a significant interaction between the integration rule and contextual certainty (2-ANOVA, F1 = 5.4741, p = 0.0346). When inputs are ambiguous and the context is highly certain (context vector with values $>$ 0.99 for the target class), we find that the additive and multiplicative (Hadamard) integration rules perform similarly. (pairwise t-test, t4 = 0.1708, p = 0.8678). For lower levels of certainty (context vector with values <=<=< = 0.6 for the target class), we note the following observations: 1) when context is matching, Hadamard integration is outperformed by Sum integration (pairwise t-test, t4 = -4.7677, p = 0.0089, and 2) when context is contradictory, Hadamard integration outperforms Sum integration (pairwise t-test, t4 = 8.9341, p = 0.0009; Fig. 3b).
@@
Figure 3: Role of integration rules on ambiguity resolution. a, comparison between Hadamard and additive integration rules. b, test set accuracy relative to contextual certainty using either integration rule in conditions where inputs are ambiguous and context is relevant (ambig.Match) or where inputs are unambiguous and context is contradictory to sensory input (unambig.Contra).